Okay im not exactly sure what to do here.
The half-life of radon-222 is about 85 hours. What fraction of an intial amount of the material would be left afer 10 hours?
Thanks for the help!

Answer by stanbon(75887)

The half-life of radon-222 is about 85 hours. What fraction of an intial amount of the material would be left afer 10 hours?
=================
Half-life formula:
A(t) = Ao (1/2)^(t/85)
This formula says the amount of material you will have
after time t is the original amount (Ao) multiplied by (1/2) each time
the total elapsed time is 85 hours.
---------------
For t = 10 hours you get:
A(10) = Ao (1/2)^(10/85)
A(10) = Ao (1/2)^(2/17)
A(10)= Ao [0.92]
The fraction that remains is 92/100 th of the original amount.
----------
Cheers,
Stan H.